# SpaceControl: Introducing Test-Time Spatial Control to 3D Generative Modeling

### v1

@@ Abstract
Generative methods for 3D assets have recently achieved remarkable progress, yet providing intuitive and precise control over the object geometry remains a key challenge. Existing approaches predominantly rely on text or image prompts, which often fall short in geometric specificity: language can be ambiguous, and images are cumbersome to edit. In this work, we introduce SpaceControl, a training-free test-time method for explicit spatial control of 3D generation. Our approach accepts a wide range of geometric inputs, from coarse primitives to detailed meshes, and integrates seamlessly with modern pre-trained generative models without requiring any additional training. A controllable parameter lets users trade off between geometric fidelity and output realism. Extensive quantitative evaluation and user studies demonstrate that SpaceControl outperforms both training-based and optimization-based baselines in geometric faithfulness while preserving high visual quality. Finally, we present an interactive user interface that enables online editing of superquadrics for direct conversion into textured 3D assets, facilitating practical deployment in creative workflows. Find our project page at https://spacecontrol3d.github.io/.
@@ 1 Introduction
Generating 3D assets is a fundamental step in building virtual worlds, useful for gaming, simulation, virtual reality applications, and digital design. Recently the field of 3D generation gained immense traction, and we are now able to create assets of previously unseen quality (Xiang et al., 2025; Zhang et al., 2024; Vahdat et al., 2022; Gao et al., 2022; Wu et al., 2025; Siddiqui et al., 2024; Zhao et al., 2025; Chen et al., 2025b; Huang et al., 2025; Corsetti et al., 2025). A persistent challenge, however, is controllability, i.e., how users can effectively steer generation to align with desired shapes and appearances.
Current controllable 3D generation methods rely mainly on text or image conditioning. Text is accessible and flexible but inherently ambiguous and ill-suited for specifying precise geometry. Images provide stronger alignment with 3D structures but are cumbersome to edit and not intuitive for fine-grained adjustments. As a result, neither modality enables artists or designers to directly manipulate the geometry of generated objects. A more natural paradigm is to allow users to interact with the generative model in 3D space, starting from coarse or abstract geometry and refining toward detailed assets.
Existing methods that introduce 3D geometric control fall into two categories: training-based and guidance-based. Training-based methods fine-tune existing generative models to support a specific form of geometric input, e.g., LION (Vahdat et al., 2022) for voxel conditioning, and Spice-E (Sella et al., 2024) for primitive or mesh conditioning. These methods provide controllability but require retraining, which reduces the original model’s generalization capabilities. In contrast, guidance-based methods such as LatentNeRF (Metzer et al., 2023) and Coin3D (Dong et al., 2024) act solely at inference time without retraining, but usually involve substantial optimization overhead and constrain 3D structure only indirectly. Other works enrich existing 3D assets with geometric and appearance detail (Michel et al., 2022; Chen et al., 2023; Barda et al., 2025), yet they assume fine-grained input geometry, limiting usability in creative workflows where artists often begin with coarse sketches.
@@
In this work, we present SpaceControl, a training-free method that injects explicit geometric control into modern framework for text- or image-conditioned 3D generation, such as Trellis (Xiang et al., 2025) or SAM 3D (Chen et al., 2025a), by directly encoding user-specified geometry into its latent space and using it as explicit guidance. Our method requires no additional training and enables controllable generation from diverse forms of geometry, ranging from simple primitives to detailed meshes.
@@
We compare SpaceControl against both training-based (Sella et al., 2024) and guidance-based (Dong et al., 2024) approaches, as well as a stronger training-based variant of Spice-E adapted to Trellis. Remarkably, despite requiring no fine-tuning, SpaceControl achieves superior geometric faithfulness while preserving visual realism. We further provide a user interface that allows online editing of superquadrics and real-time generation of textured assets, supporting practical deployment in design workflows.
In summary, our contributions are the following:
@@
  • We introduce a training-free guidance method that conditions a powerful pre-trained generative model (Trellis) on user-defined geometry via latent space intervention, enabling geometry-aware generation without the need for costly fine-tuning.
@@
  • We conduct extensive evaluations, including a user study and quantitative analysis, showing that our method outperforms prior state-of-the-art methods for shape-conditioned 3D asset generation.
@@
  • We develop an interactive user interface that enables online editing of superquadrics and their real-time conversion into detailed, textured 3D assets, supporting practical deployment in creative workflows.
@@ 2 Related Work
The field of 3D generation has experienced a rapid growth during the past few years both in terms of output modalities and controllability. Similar to the first image diffusion models (Ramesh et al., 2021), early applications of diffusion models for 3D generation (Nichol et al., 2022) were conducting the diffusion process in the original input space and were limited in the generated output type. More recent approaches (Vahdat et al., 2022; Jun & Nichol, 2023) started running the generation in a more compact latent space, leading to substantial improvements both in terms of quality and efficiency. To achieve an even increased efficiency,  (Zhang et al., 2024; Xiang et al., 2025) have started to disentangle the modeling of the structure from the appearance, leading to unprecedented high-quality generations. The separate modeling of geometry and appearance opens the door to explicit forms of spatially grounded conditioning, as done in our SpaceControl.
@@ 2.2 Controllable Generative Models
Given a pretrained generative model, there are two main approaches to introduce a new control modality: (1) methods which finetune a part or the whole network to take new types of conditioning as input, and (2) training-free methods which condition the generation via inference-time guidance. In the last years, many approaches have been developed to control the generation of image generative models, enabling conditioning in several forms as strokes, depth maps, and human poses. In contrast, analogous mechanisms for 3D generation remain far less mature.
@@
A wide variety of methods have been proposed to introduce new control modalities to image generative models. Among works based on finetuning, we identify two main lines of research. On one side, there are works based on ControlNet (Zhang et al., 2023; Bhat et al., 2024) which add conditional control to a section of the network by introducing a trainable copy connected to the original via zero convolutions. The key idea is to learn to control the original network without throwing information from the original training. On the other side, there are approaches which add additional layers for additional control of the network (Garibi et al., 2025; Hertz et al., 2022). Among training-free methods (Von Rütte et al., 2024; Meng et al., 2022; Sajnani et al., 2025), one closely related to our work is SDEdit (Meng et al., 2022) which uses stroke paintings to condition the generation of SDE-based generative models for images, by leveraging the denoising process of SDE-based generative models.
@@
Only limited works have explored spatially grounded control of 3D generative models. On one side, approaches as LatentNERF (Metzer et al., 2023), Fantasia3D (Chen et al., 2023), Instant3dit (Barda et al., 2025), and Phidias (Wang et al., 2025) apply the spatial control on image generative models by projecting the conditioning 3D model on multiple views and then leveraging test-time optimization to combine them in a 3D representation. On the other side, Spice-E (Sella et al., 2024) directly apply the control in the 3D space, by finetuning Shap-E (Jun & Nichol, 2023) separately on chairs, tables and airplanes from ShapeNet (Chang et al., 2015). Both the approaches attempt explicit spatial control, but nonetheless fall short of introducing a method that’s as usable in unconstrained settings as introduced in their 2D counterparts. The former still requires long optimization times and use the geometric input to condition the generation of the 2D projections of the 3D objects, instead of directly conditioning in 3D. The latter needs class-specific fine-tuning which limits the applicability in unconstrained settings and does not allow to model the strength of the geometric control. By contrast, SpaceControl introduces test-time guidance directly within 3D generative models, providing a framework that is efficient, accurate, and fast.
@@ 3 Preliminaries
Before introducing our SpaceControl, we review the foundations on which it builds: rectified flow matching, the Trellis generative model, as well as superquadrics.
@@ 3.2 Trellis
Trellis (Xiang et al., 2025) is a recent 3D generative model which employs rectified flow models to generate 3D assets from either textual or image conditioning. Specifically, it consists of two separate steps of generations, where the first aims to generate the structure, while the second focus on the appearance. Here, we rely on Trellis, however the very recent SAM 3D follows the same approach.
3.2.1 Structure generation
In the first stage, a noisy latent variable $\mathbf{z}_{1}\in\mathbb{R}^{16\times 16\times 16\times 8}$ is sampled from $\mathcal{N}(\mathbf{0},I)$ and denoised by a rectified flow model iteratively applying Eq. 1 using either image or text conditioning. Specifically, text conditions are encoded via the CLIP (Radford et al., 2021) text encoder, while image conditions are encoded via DINOv2 (Oquab et al., 2024). The denoised latent $\mathbf{z}_{0}$ is then decoded by a decoder $\mathcal{D}$ into a binary occupancy grid $\mathbf{x}\in\{0,1\}^{64\times 64\times 64}$, which encodes the spatial structure of the 3D asset. Note that the pretrained decoder $\mathcal{D}$ comes with a corresponding pretrained encoder $\mathcal{E}$. While $\mathcal{E}$ is not used during Trellis inference, it is required for our guidance mechanism, as detailed later.
3.2.2 Appearance generation
In the second stage, the $L$ active voxels are augmented with point-wise noisy latent features $\mathbf{s}_{1}\in\mathbb{R}^{L\times 8}$ sampled from $\mathcal{N}(\mathbf{0},I)$, denoised by a second flow model, using either text or image conditioning. The clean latents $\mathbf{s}_{0}\in\mathbb{R}^{L\times 8}$ can then be decoded into versatile formats such as 3D gaussians (GS), radiance fields (RF), and meshes (M) via specific decoders $\mathcal{D}_{O}=\{\mathcal{D}_{GS},\mathcal{D}_{RF},\mathcal{D}_{M}\}$.
3.3 Superquadrics
@@ 4.1 Setup
To introduce spatial control in 3D generation, the user provides a geometric conditioning signal alongside a text or image prompt. Our goal is to produce 3D assets that satisfy two key desiderata:
@@ 4.2 Approach
Next, we introduce SpaceControl (see Fig. 2) and describe how it enables guided 3D asset generation by injecting spatial guidance into a pretrained Trellis model. Because our control strategy differs between the first and second stages of generation, we outline the respective procedures in Secs. 4.2.1 and 4.2.2, respectively.
4.2.1 Structure generation
@@
where $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{0},I)$. Given $\mathbf{z}_{t_{0}}$, $\mathbf{z}_{0}$ can then obtained by iteratively applying Eq. 1 starting from $t_{0}$ and employing the original Structure Flow Model. We note that this process does not require any need of architectural changes nor training. We guide the generation with additional textual prompt, which is helpful to disambiguate the semantics of the object. As in the standard setting, the denoised latent $\mathbf{z}_{0}$ is then decoded into a final geometric structure $\mathbf{x}_{0}\in\{0,1\}^{64\times 64\times 64}$ by $\mathcal{D}$.
4.2.2 Appearance generation
Given the geometric structure generated in the first stage, we then employ either text or image conditioning to guide the generation of its appearance, by first expanding the active voxels with point-wise noisy latent features and then denoising them using the Appearance Flow Model. Notice that, even if the structure generation is always conditioned on text, image conditioning can still be used in to guide the appearance generation, allowing for finer control over the visual details (see Fig. 6(a) and Appendix).
4.3 Controlling the strength of spatial control
@@ 5.1 Comparing with State-of-the-art Methods
We compare SpaceControl to state-of-the-art baselines for the task of 3D-conditioned object generation. We compare to Spice-E (Sella et al., 2024), which fine-tunes Shap-E (Jun & Nichol, 2023) to support cuboid primitives as spatial guidance for 3D object generation. Since Spice-E is based on the Shap-E model (Jun & Nichol, 2023), to allow a fairer comparison we implement its correspondent for Trellis (Xiang et al., 2025), which we will refer to as Spice-E-T. We provide more details on its implementation and training in the Appendix. Note that Spice-E provides a separate checkpoint for shape stylization, which is used to evaluate the method on mesh conditioning, as it lead to better results. In addition, we also compare to Coin3D (Dong et al., 2024), which uses the shape-guidance to first generate a single view of the desired 3D asset, then leverage a Multi-View-Diffusion model to generate consistent multiple views, and finally extract the 3D representation using a volumetric-based score distillation sampling.
@@
To evaluate how different approaches handle geometric conditioning, we create a dataset of objects which contains the original mesh, a decomposition of it into geometric primitives, and a textual description of the asset. We use the mesh to evaluate methods on mesh-conditioned generation and geometric primitives to evaluate on shape-conditioned generation. Moreover, to evaluate both generation and generalization capabilities, we use objects of two ShapeNet (Chang et al., 2015) categories (chairs and tables) that Spice-E was explicitly trained on together with objects from the Toys4K (Stojanov et al., 2021) dataset, unseen by all methods during training. We use SuperDec (Fedele et al., 2025) to obtain the decomposition of the 3D assets into superquadrics and Gemini on rendered views to obtain a textual description of the assets from ShapeNet (Chang et al., 2015). For objects from Toys4k we use the textual description from  Xiang et al. (2025).
@@
Table 1: Comparison with Baselines. The evaluation metrics are L2 Chamfer Distance (CD) and Fréchet Inception Distance (FID). CD quantifies alignment with spatial control, while FID assesses realism. Results for SpaceControl are reported at $\tau_{0}=6$. CD scores are multiplied by $10^{3}$. indicates methods fine-tuned on chair and table. Trellis (Xiang et al., 2025) (model: txt-DiT-XL) does not offer spatial guidance, and is shown for reference only.

### v2

@@ Abstract
Generative methods for 3D assets have recently achieved remarkable progress, yet providing intuitive and precise control over the object geometry remains a key challenge. Existing approaches predominantly rely on text or image prompts, which often fall short in geometric specificity: language can be ambiguous, and images are difficult to manipulate. In this work, we introduce SpaceControl, a training-free test-time method for explicit spatial control of 3D asset generation. Our approach accepts a wide range of geometric inputs, from coarse primitives to detailed meshes, and integrates seamlessly with modern generative models without requiring any additional training. A control parameter lets users trade off between geometric fidelity and output realism. Extensive quantitative evaluation and user studies demonstrate that SpaceControl outperforms both training-based and optimization-based baselines in geometric faithfulness while preserving high visual quality. Finally, we present an interactive interface for real-time superquadric editing and direct 3D asset generation, enabling seamless use in creative workflows. Project page: https://spacecontrol3d.github.io/.
@@ 1 Introduction
Generating 3D assets is a fundamental step in building virtual worlds, useful for gaming, simulation, virtual reality applications, and digital design. Recently the field of 3D object generation gained immense traction, and we are now able to create assets of previously unseen quality (Xiang et al., 2025; Zhang et al., 2024; Vahdat et al., 2022; Gao et al., 2022; Wu et al., 2025; Siddiqui et al., 2024; Zhao et al., 2025; Chen et al., 2025b; Huang et al., 2025; Deb Sarkar et al., 2025). A continuing difficulty, however, is controllability, i.e., the ability for users to reliably guide the object generation process so that it matches the intended shapes and geometric characteristics.
Most controllable 3D generation approaches are conditioned on either text or images. Text is flexible and easy to use, but its ambiguity makes it poorly suited for specifying precise geometry. Images better constrain 3D structure, yet they are difficult to edit and unintuitive for fine-grained control. Consequently, neither modality allows artists or designers to directly manipulate the object geometry. The aim of this work is to lower the barrier to 3D asset creation by shifting control into 3D space itself. Rather than relying on text or image conditioning, this work introduces what we call spatial control: the use of 3D geometry as three-dimensional sketches that steer the synthesis of detailed 3D assets.
Existing approaches that provide spatially controlled 3D object generation fall into two categories: training-based and guidance-based methods. The former extend generative models by fine-tuning them to accept particular forms of geometric inputs, for example, voxel conditioning in LION (Vahdat et al., 2022) or primitive- and mesh-based inputs in Spice-E (Sella et al., 2024). Such approaches deliver explicit spatial control while maintaining the inference time of the base model, but require additional training, which often reduces their generalization capabilities. By contrast, guidance-based methods such as LatentNeRF (Metzer et al., 2023) operate solely at inference time, avoiding retraining, but typically incur substantial optimization overhead. Other approaches combine fine-tuning with test-time optimization. For instance, Coin3D (Dong et al., 2024) fine-tunes a multi-view generative model to produce geometry-consistent images, which are then used jointly with the geometric conditioning to optimize a geometry-aware radiance field.
Another line of works focuses on enriching existing 3D assets with geometric and appearance details (Michel et al., 2022; Chen et al., 2023; Barda et al., 2025), but typically assumes fine-grained input geometry, limiting its usefulness in creative workflows where artists tend to begin with coarse sketches.
@@
In this work, we present SpaceControl, a training-free method that introduces explicit spatial control into modern generative frameworks for text- or image-conditioned 3D generation, such as Trellis (Xiang et al., 2025) or SAM 3D (Chen et al., 2025a). The approach directly encodes user-specified geometry into the latent space as explicit guidance, requiring no additional training and enabling controllable generation from a broad range of geometric inputs, from simple geometric primitives to detailed meshes.
@@
We evaluate SpaceControl against existing training-based (Sella et al., 2024) and guidance-based (Dong et al., 2024) approaches, as well as a stronger training-based variant of Spice-E adapted for Trellis. Despite requiring no fine-tuning, SpaceControl achieves superior geometric faithfulness while maintaining high visual realism. Furthermore, we introduce a user interface for interactive superquadric editing and real-time textured asset generation, facilitating seamless integration into practical design flows.
In summary, our contributions are as follows:
@@
  • We introduce a training-free guidance method that conditions a powerful pre-trained generative model (Xiang et al., 2025) on user-defined geometry via latent-space intervention, enabling geometry-aware generation without costly fine-tuning.
@@
  • We provide extensive evaluations, including quantitative analysis and a user study, demonstrating that our method outperforms prior state-of-the-art approaches for shape-conditioned 3D asset generation.
@@
  • We develop an interactive interface for intuitive superquadric editing and real-time conversion into detailed, textured 3D assets, facilitating practical use in creative workflows.
@@ 2 Related Work
The field of 3D generation has experienced a rapid growth during the past few years both in terms of output modalities and controllability. Similar to the first image diffusion models (Ramesh et al., 2021), early applications of diffusion models for 3D generation (Nichol et al., 2022) were conducting the diffusion process in the original input space and were limited in the generated output type. More recent approaches (Vahdat et al., 2022; Jun & Nichol, 2023) started running the generation in a more compact latent space, leading to substantial improvements both in terms of quality and efficiency. To achieve greater efficiency,  Zhang et al. (2024); Xiang et al. (2025) have begun to disentangle geometric structure from its appearance, leading to unprecedented quality in 3D generations. This separation enables the isolated modification of geometry via spatial conditioning, which we introduce in this work.
@@ 2.2 Controllable Generative Models
Given a pretrained generative model, there are two main approaches to introduce a new control modality: (1) methods which finetune a part or the whole network to take new types of conditioning as input, and (2) training-free methods which condition the generation via inference-time guidance. In the last years, many approaches have been developed to control image generative models (Meng et al., 2022; Zhang et al., 2023), enabling conditioning in several forms such as strokes or depth maps. In contrast, analogous mechanisms for 3D object generation remain far less mature.
@@
A wide variety of methods have been proposed to introduce new control modalities to image generative models. Among works based on fine-tuning, we identify two main lines of research. One category follows ControlNet (Zhang et al., 2023; Bhat et al., 2024), which adds conditional control by introducing a trainable copy of the network connected via zero convolutions. The key idea is to learn control without discarding information from the original pre-trained model. Alternatively, other approaches add specialized layers to provide additional network control (Garibi et al., 2025; Hertz et al., 2022). Among training-free methods (Von Rütte et al., 2024; Meng et al., 2022; Sajnani et al., 2025), SDEdit (Meng et al., 2022) is closely related to our work. It leverages the denoising process of SDE-based generative models by using for example stroke paintings to initialize and condition image generation.
@@
Only a few works have explored spatially grounded control of 3D generative models. One category of approaches, including Latent-NeRF (Metzer et al., 2023), Coin3D (Dong et al., 2024), Fantasia3D (Chen et al., 2023), Instant3dit (Barda et al., 2025), and Phidias (Wang et al., 2025), applies spatial control to image generative models by projecting the 3D conditioning onto multiple views and leveraging test-time optimization to synthesize a 3D representation. Alternatively, Spice-E (Sella et al., 2024) applies control directly in 3D space by fine-tuning Shap-E (Jun & Nichol, 2023) on specific ShapeNet (Chang et al., 2015) categories. P2P-Bridge (Vogel et al., 2024) generates clean scene point clouds from noisy reconstructions. While both strategies attempt explicit spatial control, they fall short of the flexibility seen in 2D counterparts. The former requires lengthy optimization and conditions 2D projections rather than the 3D volume directly. The latter necessitates class-specific fine-tuning, which limits generalization and prevents tuning the strength of the geometric control. In contrast, SpaceControl introduces test-time guidance directly within 3D generative models, providing a framework that is efficient, accurate, and fast.
@@ 3 Preliminaries
Before introducing our SpaceControl, we review the foundations on which it builds: rectified flow matching, the Trellis (Xiang et al., 2025) generative model, as well as superquadrics.
@@ 3.2 Trellis
Trellis (Xiang et al., 2025) is a recent 3D generative model which employs rectified flow models to generate 3D assets from either textual or image conditioning. Specifically, it consists of two separate steps of generations, where the first aims to generate the geometric structure, while the second focus on the appearance. Note that SpaceControl is equally applicable to more recent works such as SAM 3D (Chen et al., 2025a), which follow the same two-step generation approach.
Structure Generation.
In the first stage, a noisy latent variable $\mathbf{z}_{1}\in\mathbb{R}^{16\times 16\times 16\times 8}$ is sampled from $\mathcal{N}(\mathbf{0},I)$ and denoised by a rectified flow model iteratively applying Eq. 1 using either image or text conditioning. Specifically, text conditions are encoded via the CLIP (Radford et al., 2021) text encoder, while image conditions are encoded via DINOv2 (Oquab et al., 2024). The denoised latent $\mathbf{z}_{0}$ is then decoded by a decoder $\mathcal{D}$ into a binary occupancy grid $\mathbf{x}\in\{0,1\}^{64\times 64\times 64}$, which represents the coarse spatial structure of the generated 3D asset. Note that the pretrained decoder $\mathcal{D}$ is accompanied by a corresponding pretrained encoder $\mathcal{E}$. While $\mathcal{E}$ is not utilized during standard Trellis inference, it is essential for our guidance mechanism, as it maps the spatial conditioning signal into the shared latent space (see Sec. 4).
Appearance Generation.
In the second stage, the $L$ active voxels of the binary occupancy grid are augmented with point-wise noisy latent features $\mathbf{s}_{1}\in\mathbb{R}^{L\times 8}$ sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$, which are denoised by a second flow model using either text or image conditioning. The clean latents $\mathbf{s}_{0}\in\mathbb{R}^{L\times 8}$ can then be decoded into versatile formats such as 3D Gaussians (GS), Radiance Fields (RF), and Meshes (M) via specific decoders $\mathcal{D}_{O}=\{\mathcal{D}_{GS},\mathcal{D}_{RF},\mathcal{D}_{M}\}$. We refer the reader to Xiang et al. (2025) for further details.
Figure 3: Superquadric primitives.
Superquadrics.
@@ 4.1 Setup
To introduce spatial control in 3D generation, the user provides a geometric conditioning signal alongside a text or image prompt. Our goal is to produce 3D assets that satisfy two key criteria:
@@ 4.2 Approach
Next, we introduce SpaceControl (see Fig. 2) and describe how it enables guided 3D asset generation by injecting spatial guidance into a pretrained Trellis model. Because our control strategy differs between the first and second stages of generation, we outline each procedure in Secs. 4.2 and 4.2, respectively.
Structure generation.
@@
where $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{0},I)$. Given $\mathbf{z}_{t_{0}}$, $\mathbf{z}_{0}$ can then be obtained by iteratively applying Eq. 1 starting from $t_{0}$ and employing the original Structure Flow Model. We note that this process does not require any need of architectural changes nor training. We guide the generation with additional textual prompt, which is helpful to disambiguate the semantics of the object. As in the standard setting, the denoised latent $\mathbf{z}_{0}$ is then decoded into a final geometric structure $\mathbf{x}_{0}\in\{0,1\}^{64\times 64\times 64}$ by $\mathcal{D}$.
Appearance generation.
Given the geometric structure generated in the first stage, we then employ either text or image conditioning to guide the generation of its appearance, by first expanding the active voxels with point-wise noisy latent features and then denoising them using the Appearance Flow Model. Notice that, even if the structure generation is always conditioned on text, image conditioning can still be used to guide the appearance generation, allowing for finer control over the visual details (see Fig. 7(a) and Appendix).
Controlling the strength of spatial control.
@@ 5.1 Comparing with State-of-the-art Methods
We compare SpaceControl to state-of-the-art baselines for the task of 3D-conditioned object generation. We evaluate against Spice-E (Sella et al., 2024), which fine-tunes Shap-E (Jun & Nichol, 2023) to support cuboid primitives as spatial guidance. To ensure a fairer comparison between backbones, we implement a corresponding version for Trellis (Xiang et al., 2025), which we refer to as Spice-E-T; details on its implementation and training are provided in the Appendix. We note that Spice-E utilizes a separate checkpoint for shape stylization when evaluating mesh conditioning, as it yields superior results. Additionally, we compare to Coin3D (Dong et al., 2024), which leverages shape guidance to first generate a single view of the desired 3D asset, then uses a fine-tuned multi-view diffusion model to generate consistent views, and finally extracts the 3D representation via volumetric-based score distillation sampling.
@@
To evaluate how various approaches handle geometric conditioning, we construct a dataset comprising original meshes, their respective geometric primitive decompositions, and corresponding textual descriptions. This dataset allows for a two-fold evaluation: assessing methods on mesh-conditioned generation using the original assets and on shape-conditioned generation via the primitives. To measure both generation and generalization performance, our evaluation spans two ShapeNet (Chang et al., 2015) categories (chairs and tables) included in the Spice-E training set, alongside objects from the Toys4K (Stojanov et al., 2021) dataset, which remain unseen by all methods during training. We utilize SuperDec (Fedele et al., 2025) to decompose the 3D assets into superquadrics and employ Gemini on rendered views to generate textual descriptions for the ShapeNet assets. For the Toys4K objects, we adopt the textual descriptions provided by Trellis.
@@
Table 1: Comparison with Baselines. We evaluate faithfulness to spatial and textual control via Chamfer Distance (CD, $\times 10^{3}$) and CLIP-I, and realism via FID (texture) and P-FID (geometry). Results for SpaceControl use $\tau_{0}=6$. indicates methods fine-tuned on chair and table categories. Trellis (Xiang et al., 2025) (txt-DiT-XL) is included for reference only as it does not support spatial guidance.